# PlaneRecTR++: Unified Query Learning for Joint 3D Planar Reconstruction and Pose Estimation

### v3

@@ 7 Sparse View Experiments
Datasets.  We adopt the setup of various existing works [7, 9] to ensure a fair comparison to published baselines, and evaluate on two large-scale sparse view datasets for simultaneous planar reconstruction and camera pose estimation.
ScanNetv2 Dataset We use a more challenging sparse view split of the ScanNetv2 dataset created by [28, 53, 13, 9], consisting of 17,237/4,051 image pairs from 1,210/303 non-overlapping scenes for training/testing. Different to ScanNetv1 for monocular plane recovery, ScanNetv2 was proposed to contain a lower overlapping ratio between frames and more complex camera rotation distributions. The average overlap ratio of adjacent frames were 20.6% and 18.6% in the training and test sets, respectively. The image size is 640 × 480 for all baselines.
MatterPort3D Dataset In contrast to ScanNetv2, the sparse view version of the MatterPort3D dataset[54] is created in a semi-simulated manner, whose RGB images are rendered from approximated 3D planar meshes during the generation process of plane annotations [7], thereby mitigating real-world impact like lighting variation. However, these rendered RGB images consist solely of planes along with numerous small erroneous facets caused by approximated 3D planar meshes, increasing the level of complexity for accurate plane detection. The training set and testing set consist of 31,932 and 7,996 image pairs, respectively, exhibiting an overlap ratio of about 21.0% [9]. The image size is also kept to 640 × 480.
@@ 7.1 Baselines and their Variants
3D planar reconstruction from sparse views demand the resolution of multiple sub-tasks, encompassing relative camera pose estimation, monocular planes recovery and multi-view plane matching. We conducted a comparative analysis between representative baselines in terms of both overall results and intermediate outcomes.
@@
Public Baselines  We compared our method with several state-of-the-art and leading sparse views planar reconstruction methods, including SparsePlanes [7], PlaneFormers [8], and NOPE-SAC [9] which are all multi-stage approaches. Specifically, all these baselines begin by estimating one or more initial camera poses from learned dense features of two input frames, and subsequently refine these camera poses using predicted planes via post optimization algorithms[7], SIFT-like keypoints [7], or separate network models[8, 9]. Therefore, their primary contribution usually lies on elaborated pose refinement module while treating monocular plane prediction and initial camera pose prediction as given knowledge from external and distinct modules. These facts are also our main differences by inferring two-view planar reconstruction in a purely end-to-end manner from images without any external priors and matching supervision.
@@
As to the evaluation of camera pose estimation, we additionally consider two popular baselines SuperGlue [31] and Pose Vit [25]. Superglue is a competitive graph-based point matching network that leverages the estimated essential matrix with a RANSAC solver to obtain the relative camera pose. Therefore, Superglue inherits the intrinsic scale ambiguity of the Essential Matrix, so that we solely focus on comparing the rotation error during its evaluation. Pose Vit uses embeddings of tokenized image patches extracted from a ViT [55] and follows the principle of the Eight-Point Algorithm [56] via another ViT module to directly estimate relative camera pose between two input images. Our proposed PlaneRecTR++ draws inspiration from its philosophy but brings several important modifications to better fit unified query learning as well as end task. Specifically, we choose to use plane-aware embeddings instead of those from raw image patches, even without requiring Quadratic Position Encodings in [25]. Our method also introduces distinct differences in the attention structure, enhancing interpretability and yielding better performance.
Improved Baseline Variants  The monocular plane predictor of SparsePlanes and PlaneFormers are based on PlaneRCNN[13], while Nope-SAC utilizes an improved version of the leading PlaneTR [15] (denoted as PlaneTRP) to recover monocular planes. Motivated by [9], to make fair comparisons, we include improved baseline variants by replacing the PlaneRCNN backbone of SparsePlanes and PlaneFormers with a more powerful PlaneTRP module, termed as SparsePlanes-TRP and PlaneFormers-TRP.
NOPE-SAC and its Variants  The NOPE-SAC framework comprises four distinct modules, namely (1) a Monocular Plane Predictor (PlaneTRP), (2) a network for pixel-level pose initialization, (3) a differentiable plane matching module based on optimal transport, and (4) a network for plane-level pose refinement. To provide a more comprehensive comparison in pose estimation to the leading NOPE-SAC [9], we further introduce its two variants to better isolate the contribution of key modules.
NOPE-SAC Init. refers to the camera pose initialization network to kick off its overall procedure, i.e., module (2) above, which is solely rely on dense pixel features. We present its pose accuracy to indicate the quality of learned pose prior [9], acting as a similar role to external pose predictor in [7, 8].
NOPE-SAC Ref. refers solely to the camera pose refinement module of NOPE-SAC that produces the final camera pose given initial pose and plane correspondences.It eliminates the pose initialization network by cutting off pose related geometric score function during plane matching, while keeping the appearance score function intact. Despite retaining a multi-stage approach, NOPE-SAC Ref. shares consensus with ours in that it relies solely on monocular plane features for plane matching and then regresses camera pose.
In contrast to NOPE-SAC and other existing baselines, one thing worth to highlight is that our unified plane embedding demonstrates multi-view consistency and implicitly accomplishes plane matching, without any external supervision for plane matching nor the need for initial pose assistance. In Section 7.7, we compare our method with optimal transport (OT) within NOPE-SAC[9] requiring external priors.
@@ VI $\downarrow$ RI $\uparrow$ SC $\uparrow$ VI $\downarrow$ RI $\uparrow$ SC $\uparrow$
@0.10 m @0.60 m @$5^{\circ}$$30^{\circ}$ Normal (°) Offset (mm)
@@
7.2 Implementation Detail
7.3 Evaluation of Monocular Planes
@@ 7.4 Relative Camera Pose Evaluation
Quantitative Results. As presented in Table VII, our PlaneRecTR++ demonstrates superior performance in estimating camera pose across all metrics on the realistic ScanNetv2 dataset compared to other methods. On the MatterPort3D dataset, our approach achieves overall comparable results to the state-of-the-art Nope-SAC method, exhibiting better translation accuracy while slightly lagging behind in rotation estimation. This discrepancy can be attributed to the characteristics of input RGB images of MatterPort3D, which contains numerous erroneous tiny planes that challenge plane detection but enhance cross-view photometric consistency during rendering (Section 7), leading to a similar accuracy trend for ours on two datasets like above mentioned monocular predictions in Section 7.3.
We have found that the multi-stage design of Nope-SAC [9] enables a robust pose estimation from both image-level and plane-level information, especially when it is difficult to conduct plane recognition and matching, whereas our end-to-end pipeline relies more heavily on plane-level embeddings and thus yields slightly lower precision estimates under such extreme scenarios. The difference observed in rotation error statistics reflects this phenomenon, with our median value being smaller than that of Nope-SAC, indicating smaller typical prediction errors and higher precision for most of the test set; meanwhile, the mean value indicates the existence of large outlier error values.
NOPE-SAC Init. and NOPE-SAC Ref., relying solely on image-level features and plane-level features, respectively, exhibit much poorer performance compared to our plane-based method. This also indicates that NOPE-SAC is highly dependent on external initial priors.
@@
Qualitative Results. Figure 10 presents the visualization of the 3D plane reconstruction results achieved by PlaneRecTR++ on the ScanNetv2 [28, 53] and the MatterPort3D [54] datasets. Note that our model implicitly learns plane matching during pose inference, and we further extract and process the probability distributions of plane correspondences from our model (see 3rd column). In Figure 11, a comparison between our proposed PlaneRecTR++ and the current state-of-the-art NOPE-SAC [9] is further provided on both datasets, demonstrating that our method yields more precise plane reconstructions and recovers more accurate relative camera poses from sparse views. Even in more challenging scenarios characterized by inconsistent brightness (column 1,3), confusion caused by symmetrical repetitive patterns (column 6) and so on, our method can implicitly acquire robust correspondences for reconstruction and pose recovery, which is outperforms NOPE-SAC that relies on initial pose prior and matching supervision.
(a) QKNum.=1
(b) QKNum.=4
@@ 7.6 Ablation Studies of Model Designs
We conducted extensive ablation studies to investigate the contributions of each design choices, particularly within our plane aware cross attention layer in Section 5.1. We focus on experimenting with following two aspects: (1) cross embedding structure (CE) within the bilinear attention mechanism, and (2) the specialized design of query, key and value subdivision.
Cross Embeddings.  Our method differs from Pose ViT [25], which also employs bilinear attention, in that we cross-place plane embeddings of different input images on both sides of the bilinear attention matrix, while Pose ViT place visual features and positional encodings of the same input image, which experimentally yields better results as shown in [25].We consider that our cross plane embeddings placement follows a more intuitive discipline and better performance. To validate our idea, we follow [25] and shift our plane aware cross attention layer’s structure to the same embedding placement strategy. The experimental findings (rows 1, 2 of Table IX and X) witness a significant degradation in performance without proposed cross embeddings set-up.
@@
Moreoever, in Table IX and X, when the key and query are guaranteed to be complete, the accuracy of all metrics with VNum. $=4$ still surpasses that with VNum. $=1$ (rows 1, 4), indicating a positive contribution from the partition of value term.
On the whole, we have validated that our design not only retain the advantages of multi-head attention in standard Transformer, but also effectively capturing the distribution of plane correspondence and further enhancing model performance.
@@
7.7 Studies of Unified Plane Embedding
Consistent Planar Attributes across Frames.  After the initial single view training and following comprehensive sparse view training, rows 2,3 of Table V and Table 7.1 exhibits comparable performance in monocular plane detection on two datasets. In the rows 3,4 of Table XI, the former relies solely on similar appearance features from a single view and achieves poor matching results, whereas the latter computes a reasonable and precise plane correspondence. In Figure 13, despite of only being trained on input sparse views, our unified query embeddings exhibit promising consistency across more frames without the need for ground truth correspondence supervision.
[…]
Implicit Plane Matching.  Compared with the differentiable optimal transport (OT) [31] method from NOPE-SAC [9], our approach owns the following advantages: 1) Most importantly, we skip the requirment of pose initialization in all previous methods [7, 8, 9]. This means that there is no need for us to convert plane parameters into the same coordinate system before effectively utilize planar geometry for matching. We believe that this is a crucial factor for constructing a single-stage method. 2) We do not require explicit supervision using ground truth correspondences. Instead, only through pose supervision, our carefully designed simple network structure actively learns multi-view consistency for plane embedding. In a single forward pass, our method implicitly performs plane matching and probabilistically synthesizes pairwise plane features for pose prediction. It does not explicitly perform plane matching and input hard plane pairs to a pose refinement network [7, 9]. 3) The correspondence attention matrix formed by our network can be processed using a simple MNN (maximum nearest neighbor) operation to obtain a hard plane assignment matrix. The accuracy of this assignment matrix is comparable or even higher than previous methods with supervisions, as shown in Tabel XI
@@ 8 Conclusion
Limitations and Future work.  In spite of the state-of-the-art performance, there are still some limitations in PlaneRecTR++. For example, the rotational accuracy on scenes from Matterport3D still suffers from large outliers, which could be partly addressed using a robust kernal function on top of existing objective functions. In addition, it still remains unexplored how to extend PlaneRecTR++ to efficiently process sequential images for a larger scale of plane recontruction in an online fashion. An exciting future venue would be to leverage the learned correspondences for long-term plane-level tracking, possibly drawing recent advancements in Transformer-based video panoptic segmentation.

### v4

@@ 7 Sparse View Experiments
Datasets.  We adopt the setup of various existing works [7, 9] to ensure a fair comparison, which includes two large-scale sparse view datasets as benchmarks.
ScanNetv2 Dataset. We use a more challenging sparse view split of the ScanNetv2 dataset created by [28, 53, 13, 9], consisting of 17,237/4,051 image pairs from 1,210/303 non-overlapping scenes for training/testing. Different to ScanNetv1 for monocular plane recovery, ScanNetv2 was proposed to contain a lower overlapping ratio between frames and more complex camera rotation distributions. The average overlap ratio of adjacent frames were 20.6% and 18.6% in the training and test sets, respectively. The image size is 640 × 480 for all baselines.
MatterPort3D Dataset. In contrast to ScanNetv2, the sparse view version of the MatterPort3D dataset[54] is created in a semi-simulated manner, whose RGB images are rendered from approximated 3D planar meshes during the generation process of plane annotations [7], thereby mitigating real-world impact like lighting variation. However, these rendered RGB images consist solely of planes along with numerous small erroneous facets caused by approximated 3D planar meshes, increasing the difficulty for accurate plane detection. Therefore, evaluations on MatterPort3D may provide unfair advantages to all two-stage baselines [7, 8, 9] with a dense pixel-based pose initialization, and cause challenges to plane-based methods (like ours). We have confirmed such findings in subsequent experiments and regard MatterPort3D mainly as a complement benchmark compared to ScanNetv2 dataset. The training set and testing set consist of 31,932 and 7,996 image pairs, respectively, exhibiting an overlap ratio of about 21.0% [9]. The image size is also kept to 640 × 480.
@@ 7.1 Baselines and their Variants
3D planar reconstruction from sparse views demands the resolution of multiple sub-tasks, encompassing relative camera pose estimation, monocular planes recovery and multi-view plane matching. We conducted a comparative analysis between representative baselines in terms of both overall results and intermediate outcomes.
@@
Public Baselines  We compared our method with several state-of-the-art and leading sparse views planar reconstruction methods, including SparsePlanes [7], PlaneFormers [8], and NOPE-SAC [9] which are all multi-stage approaches. Specifically, all these baselines begin by estimating one or more initial camera poses from learned dense features of two input frames, and subsequently refine these camera poses using predicted planes via post optimization algorithms[7], SIFT-like keypoints [7], or separate network models[8, 9]. Therefore, their primary contribution usually lies in an elaborated pose refinement module while treating monocular plane prediction and initial camera pose prediction as given knowledge from external and distinct modules. These facts are also our main differences by performing two-view planar reconstruction in a purely end-to-end manner from images without any external priors and matching supervision.
@@
As to the evaluation of camera pose estimation, we additionally consider two popular baselines SuperGlue [31] and Pose ViT [25]. SuperGlue is a competitive graph-based point matching network that leverages the estimated essential matrix with a RANSAC solver to obtain the relative camera pose. Therefore, SuperGlue inherits the intrinsic scale ambiguity of the Essential Matrix, so we solely focus on comparing the rotation error during its evaluation. Pose Vit uses embeddings of tokenized image patches extracted from a ViT [55] and follows the principle of the Eight-Point Algorithm [56] via another ViT module to directly estimate relative camera pose between two input images. Our proposed PlaneRecTR++ draws inspiration from its design philosophy but brings several important modifications to better fit unified query learning as well as end task. Specifically, we choose to use plane-aware embeddings instead of those from raw image patches, even without requiring the Quadratic Position Encodings used in [25]. Our method also introduces distinct differences in the attention structure, enhancing interpretability and yielding better performance.
@@ VI $\downarrow$ RI $\uparrow$ SC $\uparrow$ VI $\downarrow$ RI $\uparrow$ SC $\uparrow$
@0.10 m @0.60 m $5^{\circ}$ $30^{\circ}$ Normal (°) Offset (mm)
@@
Improved Baseline Variants  The monocular plane predictor of SparsePlanes and PlaneFormers is based on PlaneRCNN[13], while NOPE-SAC utilizes an improved version of the leading PlaneTR [15] (denoted as PlaneTRP) to recover monocular planes. Motivated by [9], to make fair comparisons, we include improved baseline variants by replacing the PlaneRCNN backbone of SparsePlanes and PlaneFormers with a more powerful PlaneTRP module, termed SparsePlanes-TRP and PlaneFormers-TRP.
@@
NOPE-SAC and its Variants  The NOPE-SAC framework comprises four distinct modules: (1) A Monocular Plane Predictor (PlaneTRP); (2) An attention network for dense pixel-based pose initialization; (3) A supervised differentiable plane matching module based on optimal transport; (4) A neural RANSAC network for plane-level pose refinement. To provide a more comprehensive comparison in pose estimation to the leading NOPE-SAC [9], we further introduce its two variants to better isolate the contribution of key modules.
NOPE-SAC Init. refers to the camera pose initialization network to kick off its overall procedure, i.e., module (2) above, which relies solely on dense pixel features. We present its pose accuracy to indicate the quality of learned pose prior [9], acting as a similar role to external pose predictor in [7, 8].
NOPE-SAC Ref. refers to NOPE-SAC without module (2) and mainly relies on the pose refinement module for pose estimation. NOPE-SAC Ref. performs plane matching without initial pose and directly estimates relative pose from sparser views through the neural RANSAC network. Specifically, NOPE-SAC Ref. sets the initial pose to an identity matrix and cuts off initial pose related geometric score function during plane matching, while maintaining the appearance score function intact. Despite retaining a multi-stage approach, NOPE-SAC Ref. shares consensus with ours in that it relies solely on monocular plane features for plane matching and pose regression from input sparse views.
In contrast to NOPE-SAC and other existing baselines, one thing worth highlighting is that our unified plane embedding demonstrates multi-view consistency and implicitly accomplishes plane matching, without any external supervision for plane matching or the need for initial pose assistance. In Section 7.7, we compare our method with optimal transport (OT) within NOPE-SAC[9] requiring external priors.
7.2 Implementation Detail
@@ 7.4 Relative Camera Pose Evaluation
Quantitative Results.  In Table VII, our PlaneRecTR++ demonstrates superior performance in camera pose across all metrics on the realistic ScanNetv2 dataset compared to other methods. On the MatterPort3D dataset, our approach achieves overall comparable results to the leading NOPE-SAC, exhibiting better translation accuracy while slightly lagging behind in rotation estimation.
The relatively poor rotation performance can primarily be attributed to the simulated characteristics of input images on the MatterPort3D dataset, which contains numerous erroneous tiny planes that challenge plane detection of all methods (see Section 7.3) but enhance cross-view photometric consistency during rendering (see dataset introduction in Section 7). During inference on MatterPort3D, the dense pixel-based NOPE-SAC Init. and other pose initialization networks within baseline methods [7, 8, 9] could provide a superior initial rotation, transforming the input sparse views into closer views and thus lowering the difficulty of following plane-level pose predictions.
It is noteworthy that even on this challenging MatterPort3D dataset, our median rotation error remains smaller than that of NOPE-SAC, which indicates smaller typical prediction errors.
@@
Qualitative Results.  Figure 10 presents the visualization of the 3D plane reconstruction results achieved by PlaneRecTR++ on the ScanNetv2 [28, 53] and the MatterPort3D [54] datasets. Note that our model implicitly learns plane matching during pose inference, and we further extract and process the probability distributions of plane correspondences from our model (see 3rd column). Our method exhibits superior performance in plane matching and reconstruction even in the presence of extremely sparse views, and is capable of identifying and exploiting the discontinuous ground (rows 5,7,8). In Figure 11, a further comparison between our proposed PlaneRecTR++ and the current state-of-the-art NOPE-SAC [9] is provided on both datasets, demonstrating that our method yields more precise plane reconstructions and recovers more accurate relative camera poses from sparse views. Even in more challenging scenarios characterized by inconsistent brightness (column 1,3), confusion caused by symmetrical repetitive patterns (column 6) and so on, our method can implicitly acquire robust correspondences for reconstruction and pose recovery, which is outperforms NOPE-SAC that relies on initial pose prior and matching supervision.
@@
Inference Time.  We calculate the average inference time of the latest neural methods for joint planar reconstruction and pose estimation on a NVIDIA TITAN V GPU. Our single-stage PlaneRecTR++ ($0.258$ s) achieves higher inference speed than previous multi-stage NOPE-SAC ($0.274$ s) and PlaneFormers ($4.257$ s), thanks to the exemption from external pose initialization modules [7, 8, 9] and iterative refinement modules [8].
@@ 7.6 Ablation Studies of Model Designs
We conducted extensive ablation studies to investigate the contributions of each design choice, particularly within our plane aware cross attention layer in Section 5.1. We focus on experimenting with the following two aspects: (1) cross embedding structure (CE) within the bilinear attention mechanism, and (2) the specialized design of query, key and value subdivision.
Cross Embeddings.  Our method differs from Pose ViT [25], which also employs bilinear attention, in that we cross-place plane embeddings of different input images on both sides of the bilinear attention matrix, while Pose ViT places visual features and positional encodings of the same input image, which experimentally yields better results as shown in [25]. We consider that our cross plane embeddings placement follows a more intuitive discipline and better performance. To validate our idea, we follow [25] and shift our plane aware cross attention layer’s structure to the same embedding placement strategy. The experimental findings (rows 1, 2 of Table IX and X) show a significant degradation in performance without proposed cross embeddings set-up.
@@
(a) QKNum.=1
(b) QKNum.=4
Images Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Comparison of camera poses and 3D reconstruction with $\geq 3$ views on the ScanNetv2 dataset. The entire reconstruction process of $n$ views is obtained by executing n1n-1 inferences between every two adjacent views. The fixed Black frustums show the camera of the first image. The Green frustums represent the ground truth camera of the other images while the Blue frustums depict the predicted results of the leading NOPE-SAC or ours.
@@
Moreoever, in Table IX and X, when the key and query are guaranteed to be complete, the accuracy of all metrics with VNum. =4=4 still surpasses that with VNum. =1=1 (rows 1, 4), indicating a positive contribution from the partition of value term.
On the whole, we have validated that our design not only retains the advantages of multi-head attention in standard Transformer, but also effectively captures the distribution of plane correspondence and further enhancing model performance.
7.7 Studies of Unified Plane Embedding
[…]
@@ 7.7 Studies of Unified Plane Embedding
Consistent Planar Attributes across Frames.  After the initial single view training and following comprehensive sparse view training, rows 2,3 of Table V and Table 7.1 exhibit comparable performance in monocular plane detection on two datasets. In the rows 3,4 of Table XI, the former relies solely on similar appearance features from a single view and achieves poor matching results, whereas the latter computes a reasonable and precise plane correspondence. In Figure 14, despite only being trained on input sparse views, our unified query embeddings exhibit promising consistency across more frames without the need for ground truth correspondence supervision. Figure 13 exhibits our qualitative results that conspicuously outperform the leading NOPE-SAC on multiple views ($\geq 3$ views).
Implicit Plane Matching.  Compared with the differentiable optimal transport (OT) [31] method from NOPE-SAC [9], our approach owns the following advantages: (1) Most importantly, we skip the requirement of pose initialization in all previous methods [7, 8, 9]. This means that there is no need for us to convert plane parameters into the same coordinate system before effectively utilizing planar geometry for matching. We believe that this is a crucial factor for constructing a single-stage method. (2) We do not require explicit supervision using the ground truth correspondences. Instead, only through pose supervision, our carefully designed simple network structure actively learns multi-view consistency for plane embedding. In a single forward pass, our method implicitly performs plane matching and probabilistically synthesizes pairwise plane features for pose prediction. It does not explicitly perform plane matching and input hard plane pairs to a pose refinement network [7, 9]. (3) The correspondence attention matrix formed by our network can be processed using a simple MNN (maximum nearest neighbor) operation to obtain a hard plane assignment matrix. The accuracy of this assignment matrix is comparable or even higher than previous methods with supervisions, as shown in Table XI.
@@ 8 Conclusion
Limitations and Future work.  It still remains unexplored how to extend PlaneRecTR++ to efficiently process more sequential images for a larger scale of planar recontruction in an online fashion. An exciting future venue would be to leverage the learned correspondences for long-term plane-level tracking, possibly drawing recent advancements in Transformer-based video panoptic segmentation.
Acknowledgments
We thank the anonymous reviewers and editors for their valuable comments. This work was supported in part by the NSFC (No. 62325211, No. 62132021, No. 62201603, No. 62571535), the HNNSF (No. 2025JJ40060), and the CPSF (No. 2023TQ0088, No. GZC20233539).